• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 23:59:25 on April 10, 2007

    Entry number 198513

    keyword=Shift Summary (Tuesday SWING)


    Shift Summary for April 10, 1600-2400


    Shift Crew:

    Shift Leader: Brad Sawatzky
    Target Operator: Javier Gomez

    Kinematics:

    Beam Energy: 0.9914 GeV
    Left HRS: 0.59640 GeV/c @ 140.3 degrees
    Right HRS: 0.72620 GeV/c @ 14.9 degrees

    Summary:


    Time    Description
    1600    Started shift waiting for beam (MCC working on Hall B optics
    1640    Went to Controlled Access at the request of MCC
                - MCC replaced a bad card on the Compton ion chamber
                - Ed adjusted the iris on the target camera
                - Ed fixed the beam-right Hall camera
    1940    Left Q1 tripped (due to drop out in 4K flow?) -- recovery was easy
    2320    MCC is now steered up to the BSY, but now there is a problem with
            a valve in Hall B.  Hall B is going to Controlled now...
    2330    MCC spontaneously put us to Beam Permit (we've been sitting in
            Controlled Access up to now).
    2400    Handing off -- still waiting for beam.
    

    Run List:


    Run     Target      Events      Comments
    Ziltch, Zip, Nada, Nothing.